OpenStudy (anonymous):

Someone please help me set this up correctly!!!! The property restoration company PuroServ is considering switching to new dehumidifiers. Their market research, considering the cost of the new machines and their efficiency, tells them that the switch would give them an 67% chance of making a $20,000 profit, a 3% chance of breaking even, and a 30% chance of losing $5,000. How much money does PuroServ expect to make with their new purchase?

OpenStudy (anonymous):

Answer is obtained by multiplying the corresponding profit (or loss) with their respective percentages and adding (or subtracting for the loss) 67% of 20,000 + 3% of 0 - 30% of 5,000 = 0.67* 20,000 + 0 - 0.3* 5,000 =13400 - 1500 = $ 11,900

OpenStudy (anonymous):

oh wowwww I had that set up so incorrectly. thanks so much

OpenStudy (anonymous):

You're welcome my friend